how to find the equation of a circle which touches x-axis and its centre (h,k) is given..??

Grade:12

4 Answers

vikas askiitian expert
13 years ago

if circle touches x axis then always y coordinate of center is radius of that circle ..

 r = k

eq of circle : (x-h)2 + (y-k)2 = k2

                  x2 + y2 - 2xh - 2yk + h2 = 0

this is the required eq of circle....

vijesh jayan
13 years ago

(X-H)2+(Y-K)2=R2

(H,K) IS CENTRE

TOUCHES X AXIS MEANS X HAS SOME VALUE AND Y =0

THEREFORE  (X-H)2+(-K)2=R2

(X-H)2+K2 = R2

IS WHAT I THINK

APPROVE PLZZZZZZZZZZZZZZZZZZZZZZZZZZZZZ

Surbhi Kumari
13 years ago

since the centre is given u can use the center - radius form of the equation i.e.

(x-h)^2+(Y-k)^2=r^2

where r=k

Aiswarya Ram Gupta
13 years ago

thnx